3.3.35 Cans of regular soda have volumes with a mean of 13 32 oz and a standard deviation of 0 12 oz Is it unusual for a can to contain 13 68 oz of soda? Minimum usual valueoz (Type an integer or a decimal)
Answer #1

Usual value range is the range of two standard deviations. Hence,

Minimum usual value = Mean - 2*SD = 13.32 - 2*0.12 = 13.32 - 0.24 = 13.08 oz